# Attacking Voice Anonymization Systems with Augmented Feature and Speaker Identity Difference

### v1

@@ II-A Data Augmentation Enhanced Feature Representation
where $\mathbf{X}$ is the log-Mel feature of the input speech, $\mathbf{M}_{tf}$ is the time-frequency-masking matrix applied by SpecAugment, and $\odot$ denotes element-wise multiplication. $\mathcal{F}(\cdot)$ represents the ECAPA-TDNN feature extractor, which is optimized by speaker classification with additive angular margin loss [6]. For anonymization systems that retain more speaker information (e.g., B3 and B4), the contrastive loss [7] is added to further enhance the feature representation.
@@ II-B Speaker Identity Difference Enhanced Classifier
The optimization of the PLDA classifier increases the similarity score for speaker feature embeddings belongings to the same speaker while decreasing it for embeddings from different speakers, thereby emphasizing distinctions in speaker identity. By leveraging PLDA as the classifier, our system enhances speaker identity difference during verification, thereby improving the overall attack performance.
@@ III Results
Ablation Study: To further evaluate the effects of DA representation and SID-enhanced classifier, we conduct an ablation study, as shown in Table II. The results show a significant degradation in EER performance when either of these components is removed (denoted as w/o), confirming that both DA representation and SID-enhanced classifier are effective in improving attack performance. Furthermore, the results indicate that SID-enhanced classifier provides a greater performance boost compared to DA representation, highlighting it as a more impactful design choice.
Note that due to the lack of details about the anonymization strategy, DA-SID cannot be applied to the T10-2 anonymization system in the challenge. Instead, we use a pretrained model TitaNet-Large [9] to extract speaker embeddings and a cosine similarity classifier for speaker verification, which still outperforms the official baseline.
@@ IV Conclusion
We proposed a DA-SID attacker system to address the significant feature distribution differences between original and anonymized speech. By integrating data augmentation with probabilistic linear discriminant analysis, our system effectively mitigates distribution gaps and enhances speaker identity difference. Experimental results demonstrate that our system outperforms the official baseline, which was ranked among the top-5 systems in the First VoicePrivacy Attacker Challenge.

### v2

@@ II-A Data Augmentation Enhanced Feature Representation
where $\mathbf{X}$ represents the log-Mel feature of the input speech, $\mathbf{M}_{tf}$ denotes the time-frequency masking matrix applied by SpecAugment, and $\odot$ indicates element-wise multiplication. $\mathcal{F}(\cdot)$ represents the ECAPA-TDNN feature extractor, optimized using the additive angular margin loss [6]. For anonymization systems that retain more speaker information (i.e., B3 and B4), a contrastive loss is equally incorporated to enhance the feature representation, as in [7].
@@ II-B Speaker Identity Difference Enhanced Classifier
The optimization of the PLDA classifier increases the similarity score for anonymized speaker feature embeddings belongings to the same speaker while decreasing it for embeddings from different speakers, thereby emphasizing distinctions in speaker identity. By leveraging PLDA as the classifier, our system enhances speaker identity difference during verification, thereby improving the overall attack performance.
@@ III Results
Ablation Study: To further evaluate the effects of DA representation and SID-enhanced classifier, we conduct an ablation study, as shown in Table II. Here DA is optimized with additive angular margin loss. The results show a significant degradation in EER when either of these components is removed (denoted as w/o), confirming that both DA representation and SID-enhanced classifier are effective in improving attack performance. Furthermore, the results indicate that SID-enhanced classifier provides a greater performance boost compared to DA representation, highlighting it as a more impactful design.
Note that, as the distributions of the anonymized speech pairs for T10-2 anonymization system are mismatched, DA-SID cannot be applied to T10-2 in the challenge. To address this limitation, we employ a pretrained model TitaNet-Large [9] to extract speaker embeddings, combined with a cosine similarity classifier for speaker verification. Our method achieves an EER of 32.23%, notably outperforming the official T10-2 baseline system of 41.10%.
@@ IV Conclusion
We proposed a DA-SID attacker system to address the significant feature distribution differences between original and anonymized speech. By integrating data augmentation with probabilistic linear discriminant analysis, our system effectively mitigates distribution gaps and enhances speaker identity difference. Experimental results demonstrate that our system outperforms the official baseline, which was ranked among the top-5 systems in the challenge.